# Recoverable Compression: A Multimodal Vision Token Recovery Mechanism Guided by Text Information

### v2

@@ Abstract
With the advancement of large-scale language modeling techniques, large multimodal models combining visual encoders with large language models have demonstrated exceptional performance in various visual tasks. Most of the current large-scale multimodal models achieve this by mapping visual features obtained from the visual encoder into a large language model and using them as inputs alongside text for downstream tasks. Therefore, the number of visual tokens directly affects the training and inference speed of the model. There has been significant work on token pruning for visual transformers, but for large multimodal models, only relying on visual information for token pruning or compression may lead to significant loss of important information. On the other hand, the textual input in the form of a question may contain valuable information that can aid in answering the question, providing additional knowledge to the model. To address the potential oversimplification and excessive pruning that can occur with most purely visual token pruning methods, we propose a text information-guided dynamic visual token recovery mechanism that does not require training. This mechanism leverages the similarity between the question text and visual tokens to recover visually meaningful tokens with important text information while merging other less important tokens. Experimental results demonstrate that our proposed method achieves comparable performance to the original approach while compressing the visual tokens to an average of 10% of the original quantity. Our source code will be made publicly available following acceptance.
@@ Introduction
The multimodal large language models (MM-LLMs) (Liu et al. 2023a, 2024b; OpenAI 2023; Team et al. 2023) take LLMs as the core and use LLM’s powerful language generation, zero-shot transfer, and context learning capabilities to solve multi-modal tasks. Specifically, MM-LLMs use a vision encoder, such as the NFNet-F6 (Brock et al. 2021), Vision Transformer (ViT) (Dosovitskiy et al. 2021) and the CLIP (Radford et al. 2021), to convert them into vision features, align them with the text input of LLMs through the projection layer, and finally concatenate the vision features and text features into the LLMs.
@@
All experiments were conducted in the PyTorch framework on four NVIDIA 4090 24G GPUs. We utilized the LLaVA1.5-7B with Lora fine-tuning as our baseline. The visual and text encoder is CLIP with ViT as the backbone, where the input image size is 336x336 and the patch size is 14x14. It’s worth noting that we followed the CLIP pretraining approach for handling images of different resolutions, directly resizing them to 336x336. The parameter $k$ in LOF is set to $20$ in ScienceQA, TextVQA, and MMBench. In the MME, VQAv2, and POPE $k$ was set to $30$, $90$, and $30$ respectively.
@@
Secondary Recovery Mechanism
@@
As shown in Figure 5, The tokens used for visual score screening are disorganized and do not contain the image regions corresponding to the final answer. The tokens collected for text information recovery are orderly, concentrated in regions related to the question, and include the regions contained in the answer. This indicates that our proposed method can recover lost important information through textual information.
@@ Conclusion
There is still room for improvement in the performance of our method, we will integrate this method into model fine-tuning to enhance its performance further and adapt this in multiple rounds of VQA tasks in the future.

### v3

@@ Abstract
With the advancement of large-scale language modeling techniques, large multimodal models combining visual encoders with large language models have demonstrated exceptional performance in various visual tasks. Most of the current large multimodal models achieve this by mapping visual features obtained from the visual encoder into a large language model and using them as inputs alongside text for downstream tasks. Therefore, the number of visual tokens directly affects the training and inference speed of the model. There has been significant work on token pruning for visual transformers, but for large multimodal models, only relying on visual information for token pruning or compression may lead to significant loss of important information. On the other hand, the textual input in the form of a question may contain valuable information that can aid in answering the question, providing additional knowledge to the model. To address the potential oversimplification and excessive pruning that can occur with most purely visual token pruning methods, we propose a text information-guided dynamic visual token recovery mechanism that does not require training. This mechanism leverages the similarity between the question text and visual tokens to recover visually meaningful tokens with important text information while merging other less important tokens, to achieve efficient computation for large multimodal models. Experimental results demonstrate that our proposed method achieves comparable performance to the original approach while compressing the visual tokens to an average of 10% of the original quantity.
Codehttps://github.com/banjiuyufen/Recoverable-Compression
@@ Introduction
The multimodal large language models (MM-LLMs) (Liu et al. 2023a, 2024b; OpenAI 2023; Team et al. 2023) take LLMs as the core and use this powerful language generation, zero-shot transfer, and context learning capabilities to solve multi-modal tasks. Specifically, MM-LLMs use a vision encoder, such as the NFNet-F6 (Brock et al. 2021), Vision Transformer (ViT) (Dosovitskiy et al. 2021) and the CLIP (Radford et al. 2021), to convert them into vision features, align them with the text input of LLMs through the projection layer, and finally concatenate the vision features and text features into the LLMs.
@@
All experiments were conducted in the PyTorch framework on four NVIDIA 4090 24G GPUs. We utilized the LLaVA1.5-7B with LoRA fine-tuning as our baseline. The visual and text encoder is CLIP with ViT as the backbone, where the input image size is 336x336 and the patch size is 14x14. It’s worth noting that we followed the CLIP pretraining approach for handling images of different resolutions, directly resizing them to 336x336. The parameter $k$ in LOF is set to $20$ in ScienceQA, TextVQA, and MMBench. In the MME, VQAv2, and POPE $k$ was set to $30$, $90$, and $30$ respectively. In the fine-tuning experiment, we used LoRA(Hu et al. 2022) to fine-tune LLaVA1.5-7B and train one epoch with the same data.
@@
Efficiency Analysis for Secondary Recovery Mechanism
@@
Table 4: The effectiveness of the compression ratio
Ratio ScienceQA TextVQA MME
5% 67.77 (4.5%) 51.04 (4.6%) 1187.9 (4.8%)
10% 69.01 (9.7%) 55.51 (9.9%) 1284.9 (9.2%)
50% 68.52 (49.5%) 56.14 (51.1%) 1301.8 (50.9%)
Efficiency Analysis for Compression Ratio
To verify the impact of different compression ratios on performance, As shown in 4, We conducted comparisons at compression rates of approximately 5%, 10%, and 50%, The results indicate that maintaining a compression rate of around 10% can ensure both efficiency and performance simultaneously.
@@
At the same time, we conducted fine-tuning experiments under the setting of ensuring that the average number of tokens used is 10% of the original number of tokens, The experimental results indicate that our proposed method achieved further performance improvements following fine-tuning. Compared to LLaVA PruMerge, we achieved better performance on most datasets when using a similar magnitude of tokens. At the same time, our method does not require specific hyperparameter tuning for the datasets and demonstrates better generalization capabilities. Compared to CrossGET, we narrowed the performance gap following fine-tuning. When compared to the original LLaVA1.5-7B, our method reduces computational overhead by 82%, while CrossGET achieves a reduction of 31%. Furthermore, compared to LLaVA-PruMerge+, the number of tokens we utilized is half of theirs, yet we still achieved competitive performance across multiple datasets. Chat-UniVi(Jin et al. 2024) is a unified vision-language model capable of comprehending and engaging in conversations involving images and videos through a unified visual representation, which uses the strategy in ToMe(Bolya et al. 2023), also try to compress visual token in MM-LMM. Compared to it and ToMe, our method exhibits better performance. This stems from the fact that our method is tailored for large multimodal models, simultaneously considering crucial information from both the visual and text modalities.
@@
As shown in Figure 5, The tokens used for visual score screening are disorganized and do not contain the image regions corresponding to the final answer. The tokens collected for text information recovery are orderly, concentrated in regions related to the question, and include the regions contained in the answer. This indicates that our proposed method can recover lost important information through textual information. More visualization results can be found in the Appendix-1.
@@ Conclusion
There is still room for improvement in the performance of our method, we will enhance its performance further and adapt this in multiple rounds of VQA tasks in the future.
Acknowledgements
This work was supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (No.XDA0450202), Beijing Municipal Science and Technology Project (No.Z231100010323005), CAS Project for Young Scientists in Basic Research (YSBR-083), and 2035 Innovation Mission Project (No.E4J10102).
@@ Acknowledgements
OCR-related Benchmark Results
Table 7: Performance on the OCR-related Benchmark
Method DocVQA ChartQA
LLaVA1.5 22.7 (100%) 17.8 (100%)
Ours 16.7 (9.8%) 16.29 (9.9%)
To offer a more comprehensive understanding of the model’s capabilities across different scenarios. We compared our method with the base model LLaVA-1.5-7B under the training-free setting, and the experimental results are shown in Table7. Our method still has good performance in the OCR realized benchmark, with only a slight decrease in performance while significantly reducing computational consumption.